# Mixed Precision Block-Jacobi Preconditioner: Algorithms, Performance Evaluation and Feature Analysis

### v2

@@ Abstract
In this paper, we propose two mixed precision algorithms for block-Jacobi preconditioner(BJAC): a fixed low precision strategy and an adaptive precision strategy. We evaluate the performance improvement of the proposed mixed precision BJAC preconditioners combined with the preconditioned conjugate gradient(PCG) method using problems including diffusion equations and radiation hydrodynamics equations. Numerical results show that, compared to the uniform high precision PCG, the mixed precision preconditioners can achieve speedups from 1.3×1.3$\times$1.3 × to 1.8×1.8$\times$1.8 × without sacrificing accuracy. Furthermore, we observe the phenomenon of convergence delay in some test cases for the mixed precision preconditioners, and analyse the correlation between matrix features and convergence delay behaviors, some interesting conclusions are obtained which are significant and valuable for the design of more efficient mixed precision preconditioners.
Keywords. Mixed precision, block-Jacobi preconditioner, Krylov subspace method, preconditioned conjugate gradient, radiation hydrodynamics, convergence delay, matrix feature.
@@ 1 Introduction
Solving large-scale sparse linear systems arising from the discretization of partial differential equations is a crucial challenge in scientific and engineering computing. In many realistic simulations such as laser fusion, complex fluid dynamics, structural mechanics, and electronic systems, the solution of such systems has become the main performance bottleneck [1].
@@
The preconditioned Krylov subspace methods are widely used for solving sparse linear systems in practical applications [2]. The main idea of this method is to transform the original system $Ax=b$ into an equivalent preconditioned systems $M^{-1}Ax=M^{-1}b$ via constructing a preconditioner $M^{-1}$, which is an approximation of $A^{-1}$, then the preconditioned systems is solved using Krylov subspace methods. The aim of preconditioiner is to accelerate the convergence speed of Krylov subspace iterations. The typical preconditioners include the (block) Jacobi method [2], incomplete LU decomposition method [2, 3], multigrid [2, 4, 5] and domain decomposition methods [2, 6], etc. These preconditioners are widely used in practical applications and usually dominate the computation time, thus improving their efficiency is an important issue. The efficiency of preconditioners are both problem- and architecture-specific. On one hand, it is crucial to design the problem-specific algorithms to improve the convergence rate. On the other hand, performance optimization for target architecture is required to improve the floating-point operation efficiency(Flops).
@@
In recent years, mixed precision has become a new technique to improve performance. Taking advantage of low precision computation, such as less computational cost, memory storage and data movement overheads as well as lower energy consumption, mixed precision technique improves performance by employing low precision in some components of the algorithm without sacrificing the accuracy required by the application. Currently, the mixed precision preconditioner has been investigated for various numerical algorithms [7, 8, 9].
@@
The preconditioner, as mentioned above, is essentially an approximation to the inverse of the original matrix. Thus, it is a straightforward idea to design a mixed precision preconditioner using low precision. The primary advantage of low precision preconditioners is the computational cost reduction of single iteration. However, it may lead to a decrease in convergence speed or even failure in convergence. The key to achieving performance improvement is the trade-off between the computational efficiency of single iteration and the speed of convergence, which needs to be designed carefully based on the application features and the algorithm characteristics.
@@
In this paper, we consider the block-Jacobi preconditioner(BJAC), which is a kind of basic preconditioiner suitable for massively parallel computations [2, 10]. We propose two mixed precision strategies for BJAC preconditioners, fixed-mixed-precision based BJAC (fMP-BJAC) and adaptive-mixed-precision based BJAC (aMP-BJAC). Combined with the Krylov subspace iterative methods as preconditioner, the key idea of the proposed preconditioners is to use the low precision format in BJAC, while employ the high precision format in other operations in Krylov iterations. During the iterative process, the fMP-BJAC preconditioner uses a fixed low precision for both storage and computation, while the aMP-BJAC preconditioner tune its precision automatically based on the convergence behavior. The results from two test problems, include three-dimensional diffusion equations and radiation hydrodynamics equations, show the performance improvements without sacrificing accuracy. We also investigate the impact of proposed mixed precision preconditioners on the convergence behaviour of the Krylov subspace method, and find that mixed precision preconditioners may lead to convergence delays. We further analysis the application features, such as the multiscale and diagonal dominance, that associate with the convergence delays.
@@ 2 Related Work
The main idea of the current mixed precision preconditioners [11, 12, 13, 14, 15, 16, 17, 18] is to integrate low precision preconditioners into high precision (usually double precision) Krylov subspace methods. Giraud et al.(2008) [11] proposed a single precision domain decomposition preconditioner combined with a double precision CG algorithm, the results on 2D and 3D elliptic equations illustrate the improvement in reducing CPU time and storage as well as communication costs with iteration counts increasing slightly. Arioli et al.(2009) [12] theoretically and experimentally confirmed that a single precision LU preconditioner combined with a double precision FGMRES method maintains backward stability akin to double precision. Kronbichler et al.(2019) [13] investigated a single precision geometric multigrid preconditioner with a double precision CG algorithm on GPUs, demonstrating comparable discretization error to double precision while accelerating computations by 47% to 83% for the Laplace problem. Anzt et al.(2019) [14] introduced an adaptive block-Jacobi preconditioner utilizing mixed precision storage. This approach adjusts the storage precision of each diagonal block’s inverse matrix based on its condition number among double precision, single precision, half precision, and other customized low precision formats, while maintaining double precision in results. The numerical results for the SuiteSparse matrix set demonstrate that this adaptive mixed precision storage preconditioner reduces PCG algorithm solution times by 10% to 30% compared to algorithms using double precision exclusively [15]. Carson and Khan (2023) [16] analysed the sparse approximate inverses in finite precision and investigated the behaviour of the sparse approximate inverse preconditioner within five-precision GMRES-IR refinement. Chu (2023) [17] designed a mixed precision static sparse approximation inverse preconditioner and a mixed precision heuristic SPAI preconditioner on GPU. Zhang et al.(2024) [18] propose and investigate a single precision block sparse approximation inverse preconditioner on Tensor core.
The current work on mixed precision preconditioners show the potential in improving the performance for solving sparse linear systems. However, on one hand, the test problems in current work mainly arising from the model problems or the problems from SuiteSparse matrix set with small scale, on the other hand, the impact on convergence behavior of the mixed preconditioners, which is an important aspect that impacts on overall performance, is rarely be concerned in current work. Further research is required on these issues.
@@ 3.1 The Block-Jacobi Preconditioner(BJAC)
For a sparse matrix $A$, considering its block partitioning $A={(A_{ij})}_{nb\times nb}\in R^{n\times n}$, where $nb$ is the number of blocks, $A_{ij}\in R^{n_{s}\times n_{s}}$ is the sub-matrix of size $n_{s}=n/nb$. Let $D_{b}=diag\{A_{ii},i=1,2,,\dots,nb\}$ denotes the diagonal block matrix, then the block-Jacobi preconditioner(BJAC) is defined as follows [10],
$M^{-1}_{bjac}=\sum^{k-1}_{j=0}{(I-D^{-1}_{b}A)}^{j}D^{-1}_{b}$ (3.1)
@@
$\hat{D}^{-1}_{b}=\left[\begin{array}[]{ccc}\sum^{t-1}_{i=0}{(I-{D}^{-1}_{11}A_{11})}^{i}D^{-1}_{11}&&\\ &\ddots&\\ &&\sum^{t-1}_{i=0}{(I-{D}^{-1}_{nb,nb}A_{nb,nb})}^{i}D^{-1}_{nb,nb}\end{array}\right]$ (3.3)
Although the $\hat{M}^{-1}_{bjac}$ is an approximation of $M^{-1}_{bjac}$ , and the latter can be regarded as a special case of the former, for convenience, we still calling the approximation version in (3.2) the block-Jacobi preconditioner(BJAC) in this paper, and still using $M^{-1}_{bjac}$ instead of $\hat{M}^{-1}_{bjac}$ to denote approximated BJAC preconditioner in the following sections.
As shown in (3.2) and (3.3), the BJAC preconditioner has two parameters: the number of inter-block iterations $k$ and the number of intra-block iterations $t$. Then, by choosing $k$ and $t$, the preconditioners have different approximations to $A^{-1}$. In general, the larger $t$ and $k$ are, the better the BJAC approximation to $A^{-1}$ is, hence the better it accelerates the convergence of the Krylov subspace methods, meanwhile, the cost of the preconditioner is more expensive. Therefore, these two parameters need to be chosen carefully in practical problems in order to obtain the optimal performance. In particular, the BJAC preconditioner degenerates to the diagonal scaling preconditioner $D^{-1}$ if $k=t=1$.
@@ 3.2 Mixed Precision Block-Jacobi Preconditioner(MP-BJAC)
We define two precision formats: the high precision (i.e., working precision) format $p_{w}$, the low precision format $p_{l}$. Denote the low precision BJAC preconditioner in which all floating-point computations are implemented in low precision format to be $M^{-1}_{bjac}(p_{l})$. In this paper, we present two mixed precision BJAC preconditioner: the fixed low precision BJAC preconditioner(fMP-BJAC) and the adaptive precision BJAC preconditioner (aMP-BJAC).
@@ 3.2.2 Adaptive Precision BJAC Preconditioner
1
Input: $M^{-1}_{bjac},M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l},relres,adp_{tol}$
Output: $z$
2
3if  $relres\geq adp_{tol}$ then
4       $z=M^{-1}_{bjac}r$
5else
6       $z=\text{fMP-BJAC(}M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l}\text{)}$
7 end if
8return $z$
@@ 6       $z=\text{fMP-BJAC(}M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l}\text{)}$
1
Input: $M^{-1}_{bjac},M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l},relres,adp_{tol}$
Output: $z$
2
3if  $relres\geq adp_{tol}$ then
4
5      $z=\text{fMP-BJAC(}M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l}\text{)}$
6else
7       $z=M^{-1}_{bjac}r$
8 end if
9return $z$
@@ 4.1 Test Problems
Table 1: Test problems. $s$ in Diff3D-Ani(s), Diff3D-Dis(s), Diff3D-Rand(s) is the strength of the anisotropic, discontinuous and random coefficients, as described in Section 4.1.1; FDM and FVM denote Finite Difference and Finite Volume Method respectively; $N$ and $nnz$ respectively denote the size and the number of non-zero entries of the matrix.
Problem Coefficients or Modeling Discretization Mesh Size $N$ ($nnz$)
Diff3D-Const isotropic 7-point FDM $128^{3}$ 2,097,152 (14,581,760)
@@
RHD3D-1T single-temperature 7-point FVM $128^{3}$ 2,097,152 (14,581,760)
RHD3D-3T three-temperature 6,291,456 (52,133,888)
@@
$-\nabla(\kappa\nabla u)&=f\text{,\quad}x\in\Omega\\ u&=0\text{,\quad}x\in\partial\Omega\\ \end{aligned}\end{cases}$ (4.1)
@@
The radiation hydrodynamics equation (RHD) is a fundamental governing equation in the field of high energy density physics such as laser fusion and astrophysics [21, 22, 23]. The operator splitting method is commonly used in practical simulations to divide this equation into hydrodynamic and radiation-diffusion equations to be solved separately, where the radiation-diffusion equation is shown in (4.2).
@@
    Three-temperatures modeling (RHD3D-3T). The resulting matrix for discreted RHD equation in (4.2) has the following form in (4.3):
    $A=\left(\begin{matrix}A_{R}&D_{RE}&0\\ D_{ER}&A_{E}&D_{EI}\\ 0&D_{IE}&A_{I}\end{matrix}\right)$ (4.3)
    where $A_{R}$, $A_{E}$, and $A_{I}$ are discrete systems of the three temperatures equations in (4.2) with the same sparse pattern, and the matrices $D_{RE}$, $D_{RE}$, and $D_{EI}$, $D_{IE}$ are diagonal matrices reflecting the coupling of the three temperatures in (4.2).
@@
    Single-temperature modeling(RHD3D-1T): if the three-temperatures reach to equilibrium state, the equation degenerates into a scalar single-temperature equation with the matrix of $A_{R}$ in (4.3).
@@ 4.2 Test Setting
For all experiments in this paper, the BJAC and its mixed precision versions fMP-BJAC, aMP-BJAC(hl) and aMP-BJAC(lh) are all used as the preconditioner for PCG. A reduction tolerance of relative residual L2-norm(RelResNorm) $tol_{res}$ is used as the convergence criterion for PCG iteration. $tol_{res}$ is chosen depend on the problem, for the model problem Diff3D, $tol_{res}=10^{-10}$, while for the practical problem RHD3D, $tol_{res}=10^{-12}$. The zero vector is used as the initial guess solution of iterations.

### v3

@@ Abstract
In this paper, we propose two mixed precision algorithms for block-Jacobi preconditioner(BJAC): a fixed low precision strategy and an adaptive precision strategy. We evaluate the performance improvement of the proposed mixed precision BJAC preconditioners combined with the preconditioned conjugate gradient(PCG) method using problems including diffusion equations and radiation hydrodynamics equations. Numerical results show that, compared to the uniform high precision PCG, the mixed precision preconditioners can achieve speedups from 1.3×1.3$\times$1.3 × to 1.8×1.8$\times$1.8 × without lossing accuracy. Furthermore, we observe the phenomenon of convergence delay in some test cases for the mixed precision preconditioners, and analyse the correlation between matrix features and convergence delay behaviors, some interesting conclusions are obtained which are significant and valuable for the design of more efficient mixed precision preconditioners.
keywords:
Mixed Precision, Block-Jacobi Preconditioner, Krylov Subspace Method, Preconditioned Conjugate Gradient, Radiation Hydrodynamics, Convergence Delay, Matrix Feature.
@@ 1 Introduction
Solving large-scale sparse linear systems arising from the discretization of partial differential equations is a crucial challenge in scientific and engineering computing. In many realistic simulations such as laser fusion, complex fluid dynamics, structural mechanics, and electronic systems, the solution of such systems has become a main performance bottleneck [1].
@@
The preconditioned Krylov subspace methods are widely used for solving sparse linear systems in practical applications [2]. The main idea of this method is to transform the original system $Ax=b$ into an equivalent preconditioned systems $M^{-1}Ax=M^{-1}b$ via constructing a preconditioner $M^{-1}$, which is an approximation of $A^{-1}$, then the preconditioned systems is solved using Krylov subspace methods. The aim of preconditioiner is to accelerate the convergence speed of Krylov subspace iterations. The typical preconditioners include the (block) Jacobi method [2], incomplete LU decomposition method [2, 3], multigrid [2, 4, 5] and domain decomposition methods [2, 6], etc. These preconditioners are widely used in practical applications and usually dominate the computation time, thus improving their efficiency is an important issue. In practical applications, the efficiency of preconditioners are both problem- and architecture-specific. On one hand, it is crucial to design the problem-specific algorithms to improve the convergence rate. On the other hand, performance optimization for target architecture is required to improve the floating-point operation efficiency(Flops).
@@
In recent years, mixed precision has become a new technique to improve the performance of algorithms. Specifically, mixed precision technique employs low precision in some components of the algorithm without losing the accuracy required by the application, which has less computation cost, memory storage, data movement overheads and lower energy consumption. Currently, the mixed precision preconditioner has been investigated for various numerical algorithms [7, 8, 9].
@@
The preconditioner, as mentioned above, is essentially an approximation to the inverse of the original matrix. Thus, a straightforward idea is designing a mixed precision preconditioner using low precision. The primary advantage of low precision preconditioners is the computational cost reduction of single iteration. However, it may lead to a decrease in convergence speed or even failure in convergence. The key to achieving performance improvement is the trade-off between the computational efficiency of single iteration and the speed of convergence, which needs to be designed carefully based on the application features and the algorithm characteristics.
@@
In this paper, we consider the block-Jacobi(BJAC) preconditioner, which is a kind of basic preconditioiner suitable for massively parallel computations [2, 10]. We propose two mixed precision strategies for BJAC preconditioners, fixed-mixed-precision based BJAC (fMP-BJAC) and adaptive-mixed-precision based BJAC (aMP-BJAC). When applied to the Krylov subspace iterative methods, the key is using the low precision format in BJAC and employing the high precision format in other operations in Krylov iterations. During the iterative process, the fMP-BJAC preconditioner uses a fixed low precision for both storage and computation, while the aMP-BJAC preconditioner tune its precision automatically based on the convergence behavior. Experimental results of the three-dimensional diffusion equations and radiation hydrodynamics equations, show the performance improvements without sacrificing accuracy. Moreover, we also investigate that the proposed mixed precision preconditioners may lead to the convergence delays of the Krylov subspace method, which associate with the application features, such as the multiscale and diagonal dominance.
@@ 2 Related Work
For mixed precision preconditioners, a main strategy is to integrate low precision preconditioners into high precision (usually double precision) Krylov subspace methods [11, 12, 13, 14, 15, 16, 17, 18]. Giraud et al.(2008) [11] proposed a single precision domain decomposition preconditioner combined with a double precision CG algorithm, the results on 2D and 3D elliptic equations illustrate the improvement in reducing CPU time and storage as well as communication costs with iteration counts increasing slightly. Arioli et al.(2009) [12] theoretically and experimentally confirmed that a single precision LU preconditioner combined with a double precision FGMRES(Flexible Generalized Minimal Residual) method maintains backward stability akin to double precision. Kronbichler et al.(2019) [13] investigated a single precision geometric multigrid preconditioner with a double precision CG algorithm on GPUs, demonstrating comparable discretization error to double precision while accelerating computations by 47% to 83% for the Laplace problem. Anzt et al.(2019) [14] introduced an adaptive block-Jacobi preconditioner utilizing mixed precision storage. This approach adjusts the storage precision of each diagonal block’s inverse matrix based on its condition number among double precision, single precision, half precision, and other customized low precision formats, while maintaining double precision in results. The numerical results for the SuiteSparse matrix set demonstrate that this adaptive mixed precision storage preconditioner reduces PCG algorithm solution times by 10% to 30% compared to algorithms using double precision exclusively [15]. Carson and Khan (2023) [16] analysed the sparse approximate inverses in finite precision and investigated the behaviour of the sparse approximate inverse preconditioner within five-precision GMRES-IR refinement. Chu (2023) [17] designed a mixed precision static sparse approximation inverse preconditioner and a mixed precision heuristic sparse approximation inverse preconditioner on GPU. Zhang et al.(2024) [18] propose and investigate a single precision block sparse approximation inverse preconditioner on Tensor core.
The current works on mixed precision preconditioners show the potential in improving the performance for solving sparse linear systems. However, on one hand, the test problems in current work mainly arising from the model problems or the problems from SuiteSparse matrix set with small scale, on the other hand, the impact on convergence behavior of the mixed preconditioners, which is an important aspect that impacts on overall performance, is rarely be concerned in current work. Further research is required on these issues.
@@ 3.1 The Block-Jacobi Preconditioner(BJAC)
For a sparse matrix $A$, considering its block partitioning $A={(A_{ij})}_{nb\times nb}\in R^{n\times n}$, where $nb$ is the number of blocks, $A_{ij}\in R^{n_{s}\times n_{s}}$ is the sub-matrix of size $n_{s}=n/nb$. Let $D_{b}=diag\{A_{ii},i=1,2,\dots,nb\}$ denotes the diagonal block matrix, then the block-Jacobi preconditioner(BJAC) is defined as follows [10],
$M^{-1}_{bjac}=\sum^{k-1}_{j=0}{(I-D^{-1}_{b}A)}^{j}D^{-1}_{b}$ (1)
@@
$\hat{D}^{-1}_{b}=\left[\begin{array}[]{ccc}\displaystyle\sum^{t-1}_{i=0}{(I-{D}^{-1}_{11}A_{11})}^{i}D^{-1}_{11}&&\\ &\ddots&\\ &&\displaystyle\sum^{t-1}_{i=0}{(I-{D}^{-1}_{nb,nb}A_{nb,nb})}^{i}D^{-1}_{nb,nb}\end{array}\right]$
Although the $\hat{M}^{-1}_{bjac}$ is an approximation of $M^{-1}_{bjac}$ , and the latter can be regarded as a special case of the former, for convenience, we still calling the approximation version in (2) the block-Jacobi preconditioner(BJAC) in this paper, and still using $M^{-1}_{bjac}$ instead of $\hat{M}^{-1}_{bjac}$ to denote approximated BJAC preconditioner in the following sections.
As shown in (2) and (3), the BJAC preconditioner has two parameters: the number of inter-block iterations $k$ and the number of intra-block iterations $t$. Then, by choosing $k$ and $t$, the preconditioners have different approximations to $A^{-1}$. In general, the larger $t$ and $k$ are, the better the BJAC approximation to $A^{-1}$ is, hence the better it accelerates the convergence of the Krylov subspace methods, meanwhile, the cost of the preconditioner is more expensive. Therefore, these two parameters need to be chosen carefully in practical problems in order to obtain the optimal performance. In particular, the BJAC preconditioner degenerates to the diagonal scaling preconditioner $D^{-1}$ if $k=t=1$.
@@ 3.2 Mixed Precision Block-Jacobi Preconditioner(MP-BJAC)
We define two precision formats: the high precision (i.e., working precision) format $p_{w}$ and the low precision format $p_{l}$. Denote the low precision BJAC preconditioners in which all floating-point computations are implemented in low precision format to be $M^{-1}_{bjac}(p_{l})$. In this paper, we present two mixed precision BJAC preconditioner: the fixed low precision BJAC preconditioner(fMP-BJAC) and the adaptive precision BJAC preconditioner (aMP-BJAC).
@@ 3.2.1 Fixed Low Precision BJAC Preconditioner
1:$M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l}$
2:$r=\text{ConvertPrecision}(\hat{r},p_{l})$
3:$\hat{z}=M^{-1}_{bjac}(p_{l})\cdot\hat{r}$triangleright Compute in precision $p_{l}$
4:$z=\text{ConvertPrecision}(\hat{z},p_{w})$
5:return $z$
@@ 3.2.2 Adaptive Precision BJAC Preconditioner
1:$M^{-1}_{bjac},M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l},relres,adp_{tol}$
2:if $relres\geq adp_{tol}$ then
3:     $z=M^{-1}_{bjac}\cdot r$
4:else
5:     $z=\text{fMP-BJAC}(M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l})$
6:end if
7:return $z$
@@
1:$M^{-1}_{bjac},M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l},relres,adp_{tol}$
2:if $relres\geq adp_{tol}$ then
3:     $z=\text{fMP-BJAC}(M^{-1}_{bjac}(p_{l}),r,p_{w},p_{l})$
4:else
5:     $z=M^{-1}_{bjac}r$
6:end if
7:return $z$
@@ 4.1 Test Problems
Table 1: Test problems.
Problem Coefficients or Modeling Discretization Mesh Size $N$ $nnz$
Diff3D-Const isotropic 7-point FDM $128^{3}$ 2,097,152 14,581,760
@@
RHD3D-1T single-temperature 7-point FVM $128^{3}$ 2,097,152 14,581,760
RHD3D-3T three-temperature 6,291,456 52,133,888
$s$ in Diff3D-Ani(s), Diff3D-Dis(s), Diff3D-Rand(s) is the strength of the anisotropic, discontinuous and random coefficients, as described in Section 4.1.1; FDM and FVM denote Finite Difference and Finite Volume Method respectively; $N$ and $nnz$ respectively denote the size and the number of non-zero entries of the matrix.
@@
$-\nabla(\kappa\nabla u)&=f\text{,\quad}x\in\Omega\\ u&=0\text{,\quad}x\in\partial\Omega\\ \end{aligned}\end{cases}$ (4)
@@
The radiation hydrodynamics equation (RHD) is a fundamental governing equation in the field of high energy density physics such as laser fusion and astrophysics [21, 22, 23]. The operator splitting method is commonly used in practical simulations to divide this equation into hydrodynamic and radiation-diffusion equations to be solved separately, where the radiation-diffusion equation is shown in (5).
@@
    Three-temperatures modeling (RHD3D-3T). The resulting matrix for discreted RHD equation in (5) has the following form in (6):
    $A=\left(\begin{matrix}A_{R}&D_{RE}&0\\ D_{ER}&A_{E}&D_{EI}\\ 0&D_{IE}&A_{I}\end{matrix}\right)$ (6)
    where $A_{R}$, $A_{E}$, and $A_{I}$ are discrete systems of the three temperatures equations in (5) with the same sparse pattern, and the matrices $D_{RE}$, $D_{RE}$, and $D_{EI}$, $D_{IE}$ are diagonal matrices reflecting the coupling of the three temperatures in (5).
@@
    Single-temperature modeling(RHD3D-1T): if the three-temperatures reach to equilibrium state, the equation degenerates into a scalar single-temperature equation with the matrix of $A_{R}$ in (6).
@@ 4.2 Test Setting
For all experiments in this paper, the BJAC and its mixed precision versions fMP-BJAC, aMP-BJAC(hl) and aMP-BJAC(lh) are all used as the preconditioners for PCG. A reduction tolerance of relative residual L2-norm(RelResNorm) $tol_{res}$ is used as the convergence criterion for PCG iteration. $tol_{res}$ is chosen depend on the problem, for the model problem Diff3D, $tol_{res}=10^{-10}$, while for the practical problem RHD3D, $tol_{res}=10^{-12}$. The zero vector is used as the initial guess solution of iterations.